# Contextual Document Embeddings

### v1

@@ Abstract
Dense document embeddings are central to neural retrieval. The dominant paradigm is to train and construct embeddings by running encoders directly on individual documents. In this work, we argue that these embeddings, while effective, are implicitly out-of-context for targeted use cases of retrieval, and that a contextualized document embedding should take into account both the document and neighboring documents in context - analogous to contextualized word embeddings. We propose two complementary methods for contextualized document embeddings: first, an alternative contrastive learning objective that explicitly incorporates the document neighbors into the intra-batch contextual loss; second, a new contextual architecture that explicitly encodes neighbor document information into the encoded representation. Results show that both methods achieve better performance than biencoders in several settings, with differences especially pronounced out-of-domain. We achieve state-of-the-art results on the MTEB benchmark with no hard negative mining, score distillation, dataset-specific instructions, intra-GPU example-sharing, or extremely large batch sizes. Our method can be applied to improve performance on any contrastive learning dataset and any biencoder.
@@ 2 Related Work
Much research has focused on the effect of hard negatives on the performance of contrastive learning methods (Chen et al., 2020; Qu et al., 2021; Robinson et al., 2021; Wang et al., 2023). (Zhang & Stratos, 2021) observe that harder negatives provide a better approximation of the overall cross-entropy loss, but do not consider batch-level optimizations for negative selection. (Sachidananda et al., 2023) also consider contrastive batch sampling as a global optimization problem, but do not apply their technique to state-of-the-art transformer-based text embedding models. (Ma et al., 2024) use a clustering algorithm to partition a dataset into several sub-datasets, but train a different model on each sub-dataset. Our training algorithm aims to find the hardest possible batches to train text embedding model.
@@ 5 Experimental Setup
When partitioning our dataset into batches, we encode documents and queries using GTR (Ni et al., 2021) and implement our clustering algorithm on top of FAISS (Douze et al., 2024). We cluster per-domain for $100$ steps and take the best clustering out of $3$ attempts. We select NomicBERT as our pre-trained model backbone (Nussbaum et al., 2024), which has 137M parameters. We prepend all texts with short task-specific prefixes to identify each task; prefixes are listed in Section 10.7.
@@
We initialize both $M_{1}$ and $M_{2}$ using the BERT-base model from Nussbaum et al. (2024) that includes flash attention. Weights are shared between $\phi$ and $\psi$, but notably not between $M_{1}$ and $M_{2}$. For all experiments, we train with the Adam optimizer with 1000 steps of warmup to a learning rate of $2\cdot 10^{-5}$ and linearly decay to 00 throughout training. We train for three epochs unless otherwise specified. We set the maximum sequence length for all inputs to $512$ and the number of contextual inputs to $512$ (so the second-stage model has an input length of $1024$). When computing contrastive loss, we use a fixed temperature of $\tau=0.02$. When sequence dropout is enabled in our contextual architecture, we set contextual input tokens to null vectors with a uniform probability $p=0.005$. If the batch size exceeds the number of contextual documents, we randomly sample to produce contextual inputs.
@@ 7 Analysis
To confirm that the CDE model is utilizing contextual information from $\cal D$ we consider how different contextual documents help for a given docuent $d$. Figure 7 measures results on CQADupstack, a collection of Stack Exchange forum posts. We randomly sample inputs to from $\cal D$ from a domain (x-axis) and use them as input to the downstream task $d$ marked along the y-axis. We mark a square as red if its score comes within 1 point of NDCG of the top score for its domain. Generally utilizing the documents for that domain are best, but there are some cross-over interactions. Full results for each input-output task pair are shown in the appendix.
@@ 9 Acknowledgements
Thanks to Nomic and Hyperbolic for providing necessary compute required to conduct this research. This work was partially supported by Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program #2022-22072200003. JM is supported by an NSF GFRP fellowship.

### v2

@@ Abstract
Dense document embeddings are central to neural retrieval. The dominant paradigm is to train and construct embeddings by running encoders directly on individual documents. In this work, we argue that these embeddings, while effective, are implicitly out-of-context for targeted use cases of retrieval, and that a document embedding should take into account both the document and neighboring documents in context – analogous to contextualized word embeddings. We propose two complementary methods for contextualized document embeddings: first, an alternative contrastive learning objective that explicitly incorporates document neighbors into the intra-batch contextual loss; second, a new contextual architecture that explicitly encodes neighbor document information into the encoded representation. Results show that both methods achieve better performance than biencoders in several settings, with differences especially pronounced out-of-domain. We achieve state-of-the-art results on the MTEB benchmark with no hard negative mining, score distillation, dataset-specific instructions, intra-GPU example-sharing, or extremely large batch sizes. Our method can be applied to improve performance on any contrastive learning dataset and any biencoder.
@@ 2 Related Work
Much research has focused on the effect of hard negatives on the performance of contrastive learning methods Chen et al. (2020); Qu et al. (2021); Robinson et al. (2021); Wang et al. (2023). (Zhang & Stratos, 2021) observe that harder negatives provide a better approximation of the overall cross-entropy loss, but do not consider batch-level optimizations for negative selection. Hofstätter et al. (2021) cluster queries before training and show that this improves performance. Sachidananda et al. (2023) also consider contrastive batch sampling as a global optimization problem, but do not apply their technique to state-of-the-art transformer-based text embedding models. (Ma et al., 2024) use a clustering algorithm to partition a dataset into several sub-datasets, but train a different model on each sub-dataset. Our training algorithm aims to find the hardest possible batches to train text embedding model.
@@ 5 Experimental Setup
When partitioning our dataset into batches, we encode documents and queries using GTR (Ni et al., 2021) and implement our clustering algorithm on top of FAISS (Douze et al., 2024). We cluster per-domain for $100$ steps and take the best clustering out of $3$ attempts. We select NomicBERT as our pre-trained model backbone (Nussbaum et al., 2024), which has 137M parameters. We prepend all texts with short task-specific prefixes to identify each task; prefixes are listed in Section 10.7. When pooling, we pool over text tokens only, never contextual tokens.
@@
We initialize both $M_{1}$ and $M_{2}$ using the BERT-base model from Nussbaum et al. (2024) that includes flash attention. Weights are shared between $\phi$ and $\psi$, but notably not between $M_{1}$ and $M_{2}$. For all experiments, we train with the Adam optimizer with 1000 steps of warmup to a learning rate of $2\cdot 10^{-5}$ and linearly decay to 00 throughout training. For the filtering model we select nomic-embed-v1 which was trained on the same datasets (Nussbaum et al., 2024). We train for three epochs unless otherwise specified. We set the maximum sequence length for all inputs to $512$ and the number of contextual inputs to $512$ (so the second-stage model has an input length of $1024$). When computing contrastive loss, we use a fixed temperature of $\tau=0.02$. When sequence dropout is enabled in our contextual architecture, we set contextual input tokens to null vectors with a uniform probability $p=0.005$. If the batch size exceeds the number of contextual documents, we randomly sample to produce contextual inputs.
@@ 7 Analysis
To confirm that the CDE model is utilizing contextual information from $\cal D$ we consider how different contextual documents help for a given docuent $d$. Figure 7 measures results on CQADupstack, a collection of Stack Exchange forum posts. We randomly sample inputs to from $\cal D$ from a domain (x-axis) and use them as input to the downstream task $d$ marked along the y-axis. We mark a square as red if its score comes within 1 point of NDCG of the top score for its domain. Generally utilizing in-domain works best, but there are some crossover interactions.
@@ 9 Acknowledgements
Thanks to Orion Weller, Vin Sachidananda, and Zach Nussbaum for valuable feedback on this research. We would also like to acknowledge to Nomic and Hyperbolic for providing the compute necessary to conduct this research. This work was partially supported by Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program #2022-22072200003. JM is supported by an NSF GFRP fellowship.